OpenStudy (anonymous):

i need help solving this question with substitution: 2.5x+y=-2 3x+2y=0

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

yeah , i think

OpenStudy (anonymous):

can you walk me through it

OpenStudy (anonymous):

I cant really remember the steps , i learned it awhile ago...

OpenStudy (anonymous):

well this is how i learned it for instance on this problem 5x-2y=14 2x-y=5 5x-2(2x-5)=14 -2x -2x 5x-4x+10=14 -y=-2x+5 x+10=14 --- --- --- -10 -10 -1 -1 -1 x=4 y=2x-5 2(4)-y=5 8-y=5 -8 -8 -y=-3 -- -- -1 -1 y=3 ANSWER: (4,3) but i keep getting a wacky answer.

OpenStudy (anonymous):

that seems right to me .... ;/